OpenStudy (anonymous):

can some one tell me wut the square root of 75m^5n^2 is???

OpenStudy (anonymous):

\[75m^5n^2=3\times 25 \times m \times m^4 \times n^2=3m\times 5^2 \times(m^2)^2\times n^2\]